# Flow of Reasoning: Training LLMs for Divergent Problem Solving with Minimal Examples

### v4

@@ 3.2 Efficient Diversity-seeking Training of LLMs
3.2.1 Training Objective
In this work, we derive our training objective based on the trajectory balance approach (Malkin et al., 2022a), which has shown improved efficiency than other alternatives (Bengio et al., 2023; 2021). Specifically, for any complete forward trajectory $\tau=(s_{0}\to s_{1}\to\dots\to s_{n})$, the trajectory balance constraint, with a task goal $g$, says (Figure 2):
$\begin{split}Z(s_{0},g)\prod\nolimits_{t=1}^{n}P_{F}(s_{t}|s_{t-1};g)=F(s_{n})\prod\nolimits_{t=1}^{n}P_{B}(s_{t-1}|s_{t};g),\end{split}$ (4)
@@ 3.2.1 Training Objective
$\begin{split}l(\tau;\theta,g)=\left(\log\frac{Z(s_{0},g)\prod_{t=1}^{n}P_{F}(s_{t}|s_{t-1};\theta,g)}{R(s_{n})\prod_{t=1}^{n}P_{B}(s_{t-1}|s_{t};\theta,g)}\right)^{2},~{}~{}P_{B}(s_{t-1}|s_{t};\theta,g):=\frac{1}{|\text{Pa}(s_{t})|},\end{split}$ (5)
where $|\text{Pa}(s_{t})|$ denotes the number of parents of state $s_{t}$, and Malkin et al. (2022a) suggested a canonical choice of setting $P_{B}(\cdot|s_{t})$ to be uniform over the parents. Note that $Z$ is the total flow conditioning on each goal $g$ and initial state $s_{0}$. Estimating $\log Z$ can be cumbersome. We thus follow (Zhang et al., 2023a) to use the log-variance approximation, which implicitly estimates $\log Z$ given each trajectory $\tau$:
$\begin{split}\Phi(\tau;\theta)=\log R(s_{n})+\sum\nolimits_{t=1}^{n}\log P_{B}(s_{t-1}|s_{t};\theta,g)-\sum\nolimits_{t=1}^{n}\log P_{F}(s_{t}|s_{t-1};\theta,g),\end{split}$ (6)
@@
$\begin{split}\mathcal{L}_{V}(\tau;\theta)=(\Phi(\tau;\theta)-\mathbb{E}_{\tau}[\Phi(\tau;\theta)])^{2},\end{split}$ (7)
where we draw sample trajectories $\tau$ from a behavior policy $\pi(\tau;\theta,g)$ for training, and $\mathbb{E}_{\tau}[\Phi(\tau;\theta)]$ is estimated with a mini-batch of sampled trajectories. Different configurations of $\pi$ result in on-policy, off-policy, and mixed explorations, which could impact training efficiency as shown in ablation studies (§4). We discuss our method of defining $\pi(\tau;\theta,g)$ below. If $\mathcal{L}_{V}(\tau;\theta)$ is globally optimized, the resulting flow satisfies Eq.(3) and $P_{F}(\cdot|\cdot;\theta,g)$ samples proportionally to the reward as desired.
3.2.2 Efficient Exploration
Training with the loss function in Eq. (7) requires trajectory collection, raising the challenge of how to collect them efficiently and effectively. On-policy sampling, which generates trajectories via $\tau\sim P_{F}(\tau;g)$ and optimizes $\mathcal{L}_{V}(\tau;\theta)$ with a batch of trajectories, often struggles to explore the vast trajectory space due to limited exploration capabilities. This can result in inadequate estimation of the distribution needed to capture divergent reasoning paths. To address this, we want to set up a distribution $\pi(\tau,g;\theta)$ in Eq. (7) that enhances exploration while producing various training samples for policy optimization. Inspired by recent advances in GFlowNets (Vemgal et al., 2023; Hu et al., 2023a), we combine on-policy and off-policy strategies and further enhance exploration by adapting local search (Kim et al., 2023) for LLM training.
For on-policy exploration, we use both the online policy $P_{F}(s_{t}|s_{t-1};\theta,g)$ and a tempered variant to generate trajectories given a goal $g$ and initial state $s_{0}$. Off-policy exploration leverages a replay buffer prioritizing high-reward trajectories, $\epsilon$-sampling, and offline trajectory data (for Game24 in §4.3). To further explore high-reward regions, we propose a modified local search method (Figure 2). Specifically, we select the highest-reward trajectory from each batch, truncate its latter portion, and reconstruct it using a random policy $P_{U}$. This random policy avoids the computationally intensive forward process of LLMs, leading to enhanced efficiency. This approach reconstructs trajectories with a high probability of receiving higher rewards, as partially destroyed trajectories with high rewards are more likely to select the correct actions at the initial steps, while potentially making mistakes in the subsequent steps. Further details on the exploration strategies and local search process are provided in Appendices D and E, respectively.
@@ 4.1 Experimental Settings
Baselines. We compare our approach with several prompting-based methods, including $k$-shot CoT (Wei et al., 2022) (with $k$ = 1, 5, 15), ToT (Yao et al., 2024) (using BFS and DFS), GoT (Besta et al., 2024), XoT (Ding et al., 2023), and RAP (Hao et al., 2023). For fine-tuning-based methods, we evaluate SFT with diversity-enhancing decoding strategies like Temperature Sampling (Shi et al., 2024) ($\alpha=1.0,0.5,0.1$, where $\alpha$ is a temperature used to adjust the probability distribution over the vocabulary for the next token), Nucleus Sampling (Holtzman et al., 2019) ($\eta=0.95$, selecting from tokens that together make up $\eta$ of the probability mass), Typical Sampling (Meister et al., 2023) ($\tau=0.95$, where the sampling distribution is restricted to words with negative log probabilities near the conditional entropy, and then $\eta$ of the distribution mass is truncated), and diverse beam search (Vijayakumar et al., 2016) (DBS, beam width $k=20$). Additionally, we apply fine-tuning with PPO (Schulman et al., 2017) and GFN-CoT (Hu et al., 2023a). We also compare against OpenAI-O1, the latest and strongest reasoning model. All finetuning methods are trained on the same dataset as FoR. Except for Game24, which uses Llama-2-13B, Llama-3-8B is the base model for all other tasks. In the BlocksWorld task, we evaluated most baselines for a broad comparison, but some methods (e.g., GFN-CoT) showed suboptimal performance, which informed our decision to selectively apply baselines in the subsequent tasks. For the remaining tasks, we focused on high-performing methods from the initial evaluation.
Evaluation. As mentioned in §1, an effective reasoning method should not only produce correct solutions but also maximize the number of correct solutions found. Unlike previous approaches that evaluate a single solution per problem, we propose generating $n$ solutions per problem and assess methods based on four criteria: (1) Accuracy (Acc): Success is defined if at least one of the $n$ solutions is correct. (2) Diversity: For solved problems, we report the average number of unique correct solutions among the $n$—higher is better (see Appendix C.1 for details). (3) Creativity: For each method, we report the proportion of unique successful trajectories found in all solutions that are not discovered by other methods. (See Appendix C.2 for details.) (4) Runtime: The average time taken by a method to produce one solution is used as an efficiency metric. For all the datasets, we recorded the average results and standard deviation of our performance from 3 repetitions, except for tree- and graph-structured methods and O1-series, which require significant time or expense to find a single solution. Creativity is calculated based on the result of 1 repetition due to the small standard deviation observed in accuracy and diversity metrics.
4.2 Embodied Reasoning: BlocksWorld
Setup. Blocksworld examples (Valmeekam et al., 2024; Hao et al., 2023) are grouped by the minimum number of required actions: 30 examples for 2 steps, 57 for 4 steps, and 114 for 6 steps, following Hao et al. (2023). We select the first 15 of each group as the training examples for FoR and the rest as test examples. We sampled 8, 20, and 40 times for the 2, 4, and 6-step datasets, respectively, to report diversity and creativity. All the baselines are included in §4.1.
Table 1: Results on BlocksWorld, comparing prompting-based and finetuning-based methods on questions requiring two, four, and six steps, respectively. Standard deviations of three runs are shown in brackets (except for GPT-4o due to budget limit and ToT/RAP as they are exceedingly slow). We also report results from the O1-series models. Since these models are optimized for multi-step reasoning, their performance provides a reference for the upper limit of reasoning accuracy. (For O1-preview, we sampled only one solution due to budget limit.)
@@ 4.2 Embodied Reasoning: BlocksWorld
Results. As shown in Table 1, our method demonstrates improvements across all metrics. In terms of accuracy, our method outperforms the best prompting-based baseline (GPT-4o with CoT) by 80% in 4-step tasks and 16% in 6-step tasks, and exceeds the best finetuning-based baseline (SFT + Nucleus) by 82% and 86%, respectively, highlighting its robustness with increasing task complexity. From a diversity standpoint, it outperforms SFT + PPO by around 14%, showing our method is able to generate more diverse solutions. Our method outperforms all other baselines in the creativity metric, discovering more unique solutions that are not found by other baselines. Most baselines do not find any unique solutions, resulting in a creativity score of 0. It is worth noting that the O1 series improves the accuracy to a large extent, but still struggles to find divergent reasoning paths. Additionally, FoR matches the inference speed of high-efficiency baselines like $k$-shot CoT, ToT (DFS), and SFT-based methods, while being 30× faster than ToT (BFS) and RAP on a single NVIDIA A100 GPU. Training costs are detailed in Appendix C.3.
@@ 4.3 Math Puzzle Solving: Game of 24
Setup. We use the LLM-reasoner dataset (Hao et al., 2024) and randomly select 20 examples for training, with examples ranked from easy to hard and those indexed 910-1010 used for testing. Since prior works show that LLMs struggle to online sample a correct trajectory in this task (Yao et al., 2024; Yu et al., 2023a), we use Python code to generate the offline ground-truth data with diverse trajectories, which is used for fine-tuning methods. In addition, to avoid the pitfalls of arithmetic calculations with language models, we use Python code to calculate the results after "=" in an action $a_{t}$ across all methods evaluated in our experiment. We compare with baselines mentioned in §4.1 and additionally report the OpenAI-O1 mini performance. We sampled 20 times to report performance.
@@
SFT ($\alpha=1.0$) 19.00 1.37 6.49
FoR 41.00 (0.82) 1.52 (0.01) 31.82
@@
Results. As shown in Table 2, FoR demonstrates superior accuracy and diversity in solving math puzzles compared to other baselines with the same base model. Not surprisingly, O1-mini and GPT-4o achieve better performance due to the stronger intrinsic mathematical knowledge and reasoning mechanism. We also investigate the fact that GPT-4o tends to use self-verification and reflection during Game24’s inference. This may explain its superior performance in this task.
@@ 4.4 Spatial Reasoning: Rubik’s Cube
Setup. We randomly select 15 examples from the training dataset in  (Ding et al., 2023), and evaluate different methods on a test set containing 183 examples. Each example can be solved in four steps. All the baselines are included in §4.1. For SFT, CoT and FoR, 10 solutions are sampled. See more details in Appendix C.6.
Reward Design. Similar to the above tasks, a high reward is given for successful restoration. The augmented reward is based on the difference in the minimum steps required to restore from the current cube state. If an action reduces the required steps, it receives a higher reward; otherwise, it gets a lower one. Formally, $R(s_{n})=w\cdot\mathbb{I}(\text{success})+\sum_{t=1}^{n}\text{exp}(r(s_{t-1})-r(s_{t}))$, where $r(s_{t})$ represents the remaining minimum steps.
4.4.1 Results
As shown in Table 3, when there is a limited amount of training data available, our method outperforms all baselines in the Rubik’s Cube task across all three metrics. It outperforms the best baseline (XoT) by over 120% in accuracy. Additionally, our approach generates 29% more diverse solutions, while other baselines only produce one solution on average. Notably, from a creativity perspective, our approach generates a large amount of unique solutions that other baselines are unable to discover.
@@ 4.5 Abstraction Reasoning: 1D-ARC
Setup. We randomly select 5 examples from the 1d_move_1p, 1d_padded_fill, and 1d_denoising tasks. These tasks involve moving the color bar by 1 pixel, filling in empty spaces enclosed by pixels, and removing noise-like pixels, respectively. The 15 selected examples form the training set, while the remaining 45 examples from each task compose the test dataset. We sample 20 solutions for each example during inference. See Appendix C.7 for more details.
@@
Baselines. Since there are no complex reasoning baselines (e.g., ToT) evaluated on this task, we compare against Input-ouptut(IO) prompting (Xu et al., 2023b; Mirchandani et al., 2023), which involves incorporating training input-output pairs into the prompt and prompting LLMs to infer output grids given the test input directly. Program Only and Hypothesis Search (Wang et al., 2023b) synthesize Python programs for transformation, with the latter first generating language-based transformation hypotheses before program synthesis. Fine-tuning methods are not compared due to the lack of labeled reasoning data.
Reward Design. In addition to a success reward, we design the augmented rewards for actions based on how much they reduce the distance to the ground truth. Specifically, an action receives a higher reward if it reduces the hamming distance between the current state and the ground truth. The total reward is $R(s_{n})=w\cdot\mathbb{I}(\text{success})+\sum_{t=1}^{n}\sum_{i=1}^{K}\exp(h_{d}(s_{t-1}^{i},g^{i})-h_{d}(s_{t}^{i},g^{i}))$, where $h_{d}(\cdot,\cdot)$ is hamming distance and $K$ is the number of training input-output pairs. $\mathbb{I}$ indicates 1 only when the searched program successfully transforms the test input to output.
4.5.1 Results
As shown in Table 4, FoR substantially outperforms previous methods on all metrics, especially diversity and creativity. Previous approaches (hypothesis search and program-only) generate programs at once, which easily results in errors during the intermediate process, leading to inferior performance. As expected, FoR outperforms CoT, given CoT lacks mechanisms to try different solutions, causing them to rely solely on their internal knowledge and limiting their creativity.

### v6

@@ 3.2 Efficient Diversity-Seeking Training of LLMs
Training Objective. Since we can sample a complete trajectory, we derive our training objective based on the trajectory balance approach (Malkin et al., 2022a), which has shown improved efficiency than other alternatives (Bengio et al., 2023, 2021). Specifically, for any complete forward trajectory $\tau=(s_{0}\to s_{1}\to\dots\to s_{n})$, the trajectory balance constraint, with a task goal $g$, says (Figure 2):
$\small\begin{split}Z(s_{0},g)\prod\nolimits_{t=1}^{n}P_{F}(s_{t}|s_{t-1};g)=F(s_{n})\prod\nolimits_{t=1}^{n}P_{B}(s_{t-1}|s_{t};g),\end{split}$ (4)
@@
$l(\tau;\theta,g)$ $=\left(\log\frac{Z(s_{0},g)\prod_{t=1}^{n}P_{F}(s_{t}|s_{t-1};\theta,g)}{R(s_{n})\prod_{t=1}^{n}P_{B}(s_{t-1}|s_{t};\theta,g)}\right)^{2},$ (5)
$\quad P_{B}(s_{t-1}|s_{t};\theta,g):=\frac{1}{|\text{Pa}(s_{t})|}$
where $|\text{Pa}(s_{t})|$ denotes the number of parents of state $s_{t}$, and (Malkin et al., 2022a) suggested a canonical choice of setting $P_{B}(\cdot|s_{t})$ to be uniform over the parents. Note that $Z$ is the total flow conditioning on each goal $g$ and initial state $s_{0}$. The term $\log Z$ is challenging to learn accurately in an LLM setting (Hu et al., 2023a). We thus follow (Zhang et al., 2023c) to use the log-variance approximation, which implicitly estimates $\log Z$ given each trajectory $\tau$:
$\Phi(\tau;\theta)=$ $\log R(s_{n})+\sum\nolimits_{t=1}^{n}\log P_{B}(s_{t-1}|s_{t};\theta,g)$ (6)
$-\sum\nolimits_{t=1}^{n}\log P_{F}(s_{t}|s_{t-1};\theta,g),$
@@
$\small\begin{split}\mathcal{L}_{V}(\tau;\theta)=(\Phi(\tau;\theta)-\mathbb{E}_{\tau}[\Phi(\tau;\theta)])^{2},\end{split}$ (7)
where we sample trajectories $\tau$ from a behavior policy $\pi(\tau;\theta,g)$ for training, and $\mathbb{E}_{\tau}[\Phi(\tau;\theta)]$ is estimated with a mini-batch of sampled trajectories. We have different configurations of $\pi$, such as on-policy, off-policy, and mixed explorations, which could impact training efficiency (see ablation studies (§4)). If $\mathcal{L}_{V}(\tau;\theta)$ is globally optimized, the resulting flow satisfies Eq.(3) and $P_{F}(\cdot|\cdot;\theta,g)$ samples proportionally to the reward as desired.
Efficient Exploration. Training with the loss function in Eq. (7) requires trajectory collection, raising the challenge of how to collect them efficiently and effectively. On-policy sampling, which generates trajectories via $\tau\sim P_{F}(\tau;g)$ and optimizes $\mathcal{L}_{V}(\tau;\theta)$ with a batch of trajectories, often struggles to explore the vast trajectory space due to limited exploration capabilities. This can result in inadequate estimation of the distribution needed to capture diverse reasoning paths. To address this, we want to set up a distribution $\pi(\tau,g;\theta)$ in Eq. (7) that enhances exploration while producing various training samples for policy optimization. Inspired by recent advances in GFlowNets (Vemgal et al., 2023; Hu et al., 2023a), we combine on-policy and off-policy strategies and further enhance exploration by adapting local search (Kim et al., 2023) for LLM training.
For on-policy exploration, we use both the online policy $P_{F}(s_{t}|s_{t-1};\theta,g)$ and a tempered variant to generate trajectories given a goal $g$ and initial state $s_{0}$. Off-policy exploration leverages a replay buffer prioritizing high-reward trajectories, $\epsilon$-sampling, and offline trajectory data (for Game24 in §2). To efficiently explore high-reward regions, we propose a modified local search method (Figure 2). Specifically, we select the highest-reward trajectory from each batch, truncate its latter portion, and reconstruct it using a random policy $P_{U}$. This avoids the computationally expensive forward process of LLMs, enhancing efficiency. High-reward trajectories are more likely to reconstruct correctly in early steps, with errors typically occurring later. Additional details on the full algorithm, exploration strategies, and local search are in Appendices E and F.
@@ 4.1 Experimental Settings
Baselines. We first compare our approach with several prompting-based methods, including $k$-shot CoT (Wei et al., 2022), ToT (Yao et al., 2024), GoT (Besta et al., 2024), XoT (Ding et al., 2023), and RAP (Hao et al., 2023). We use the Llama-3-8B (Dubey et al., 2024) as the base model for all tasks. For fine-tuning-based methods, we evaluate SFT with diversity-enhancing decoding strategies like Temperature Sampling (Shi et al., 2024), Nucleus Sampling (Holtzman et al., 2019) ($\eta=0.95$), Typical Sampling (Meister et al., 2023) ($\tau=0.95$), and diverse beam search (Vijayakumar et al., 2016) (DBS, beam width $k=20$). Additionally, we apply fine-tuning with PPO (Schulman et al., 2017) and GFN-CoT (Hu et al., 2023a). We also compare it against OpenAI-O1, which is the latest and strongest reasoning model. All finetuning methods are trained on the same dataset as FoR. In the BlocksWorld task, we evaluated most baselines for a broad comparison, but some methods (e.g., GFN-CoT) showed suboptimal performance, which informed our decision to selectively apply high-performing baselines in the subsequent tasks.
Table 1: Results on BlocksWorld, comparing prompting-based and finetuning-based methods on questions requiring two, four, and six steps, respectively. The standard deviations of three runs are shown in brackets. $\alpha$ is a temperature. ${}^{\mathord{\clubsuit}}$ indicates that results from multiple runs for ToT (DFS).
@@
Evaluation. We generate $n$ solutions per problem and evaluate methods based on four key metrics: (1) Accuracy (Acc): A problem is considered correct if at least one of the $n$ sampled solutions is correct. (2) Diversity: For problems with at least one correct solution, we report the average number of semantically different correct solutions among the number of correct solutions, where higher is better (see Appendix for details). (3) Creativity: The proportion of unique correct solutions found by a method that are not discovered by others. (4) Runtime: The average time to generate one solution by using a single NVIDIA A100 80GB GPU is used as an efficiency metric. See Appendix C.1 and C.2 for details of metrics. [^1] In our experiments, we report the average accuracy and standard deviation over three repetitions, except for search-based methods and O1-series due to time and budget constraints. Creativity is not averaged over repetitions, as the performance exhibit low variance across runs.
4.2 Embodied Reasoning: BlocksWorld
Setup. Blocksworld examples (Valmeekam et al., 2024) are grouped by the minimum number of required actions: 30 examples for 2 steps, 57 for 4 steps, and 114 for 6 steps, following Hao et al. (2023). We select the first 15 of each group as the training examples for FoR and the rest as test examples. We sampled 8, 20, and 40 times for the 2, 4, and 6-step datasets, respectively, to report diversity and creativity. All the baselines are included in §4.1.
Results. As shown in Table 1, Our method outperforms all other baselines across all metrics, showing that FoR not only achieves the best accuracy but also finds more diverse and unique solutions. Specifically, most baselines do not find any unique solutions, resulting in a creativity score of 0. It is worth noting that the O1 series improves the accuracy to a large extent, but still struggles to find diverse reasoning paths. Additionally, FoR matches the inference speed of high-efficiency baselines like $k$-shot CoT, ToT (DFS), and SFT-based methods, while being 30× faster than ToT (BFS) and RAP. Additional results of diversity-encouraging prompts, larger-size models, and out-of-distribution (OOD) generalization are in Appendix C.4, and training costs are detailed in Appendix C.3.
@@ 4.3 Math Puzzle Solving: Game of 24
SFT ($\alpha=1.0$) 29.00 1.18 11.22
FoR 48.00 (1.41) 1.30 (0.09) 29.15
Setup. We use the LLM-reasoner dataset (Hao et al., 2024) and randomly select 20 examples for training and 100 examples for testing. Since prior work shows that LLMs struggle to sample correct trajectories online in Game24 (Yao et al., 2024; Yu et al., 2023a), we generate offline ground-truth trajectories using Python code for fine-tuning. We sample 20 solutions for all methods.
@@
Results. As shown in Table 2, FoR demonstrates superior accuracy and diversity in solving math puzzles compared to other baselines with the same base model. Not surprisingly, O1-mini and GPT-4o achieve better performance due to the stronger intrinsic mathematical knowledge and reasoning mechanism. We also investigate the fact that GPT-4o tends to use self-verification and reflection during Game24’s inference. This may explain its superior performance in Game24.
@@ 4.4 Spatial Reasoning: Rubik’s Cube
Setup. We randomly select 15 examples from the training dataset in from (Ding et al., 2023), and evaluate different methods on a test set containing 183 examples. Each example can be solved in four steps. All the baselines are included in §4.1. For SFT, CoT and FoR, 10 solutions are sampled. See more details in Appendix C.6.
Reward Design. Similar to previous tasks, we have a success reward and an augmented reward that is based on the difference in minimum steps required to restore the cube from its current state. Actions that reduce steps receive higher rewards, while others receive lower rewards. Formally, $R(s_{n})=w\cdot\mathbb{I}(\text{success})+\sum_{t=1}^{n}\text{exp}(r(s_{t-1})-r(s_{t}))$, where $r(s_{t})$ represents the remaining minimum steps.
Results. As shown in Table 3, FoR outperforms all baselines on the Rubik’s Cube task, achieving absolute increase of 5.95% in accuracy than XoT, generating an absolute 29% increase in diverse solutions, and discovering unique and creative solutions than the baselines.
@@ 4.5 Abstraction Reasoning: 1D-ARC
Setup. We randomly select 5 examples from the 1d_move_1p, 1d_padded_fill, and 1d_denoising tasks. These tasks involve moving the color bar by 1 pixel, filling in empty spaces enclosed by pixels, and removing noise-like pixels respectively. The 15 selected examples form the training set, while the remaining 45 examples from each task form the test dataset. We sample 20 solutions for each example during inference. See Appendix C.7 for details.
@@
Baselines. Since there are no complex reasoning baselines (e.g., ToT) evaluated on this task, we compare against Input-output (IO) prompting (Mirchandani et al., 2023). Program Only and Hypothesis Search (Wang et al., 2023b) synthesize Python programs for transformation, with the latter first generating language-based transformation hypotheses before program synthesis. Fine-tuning methods are not compared due to the lack of labeled reasoning data.
Reward Design. In addition to a success reward, we design the augmented rewards for actions based on how much they reduce the distance to the goal. Specifically, an action receives a higher reward if it reduces the distance between the current state and the goal. The total reward is $R(s_{n})=w\cdot\mathbb{I}(\text{success})+\sum_{t=1}^{n}\sum_{i=1}^{K}\exp(h_{d}(s_{t-1}^{i},g^{i})-h_{d}(s_{t}^{i},g^{i}))$, where $h_{d}(\cdot,\cdot)$ is hamming distance and $K$ is the number of training input-output pairs.
Method 2-step 4-step 6-step
Acc. (%) Acc. (%) Diversity Acc. (%) Diversity
FoR (Ours) 100.00 (0.00) 98.41 (1.12) 1.27 (0.02) 78.44 (4.54) 1.33 (0.03)
Results. Results in Table 4 show that FoR substantially outperforms previous methods on all metrics, especially diversity and creativity. Previous approaches (hypothesis search and program-only) generate programs at once, which easily results in errors during the intermediate process, leading to inferior performance.
Table 6: Results on GSM8K.
Method Acc. (%) Diversity-H Diversity-G
CoT (2-shot) 45.72 1.12 1.60
CoT-SC (2-shot) 41.74 - -
RAP 37.16 - -
SFT ($\alpha=1.0$) 52.69 1.13 1.63
FoR 57.39 1.26 1.72
FoR + PRM 62.62 1.31 1.77
4.6 Mathematical Reasoning: GSM8K
Setup. We use the last 50 training examples in the original training dataset, and we sample 4 times for every problem at inference. We follow the baseline implementations from LLM reasoners (Hao et al., 2024). Since there is no standard metric for reasoning diversity, we manually annotate 50 test examples to evaluate the semantic difference between solutions as the diversity metric (Diversity-H), and propose an automatic metric using GPT-4o to distinguish different solutions (Diversity-G), where the prompt is provided in Appendix C.9.
Reward Design. We removed the success reward since the pretrained mathematical knowledge enables LLMs to perform relatively well compared to BlocksWorld. Using LLMs to verify each state is computationally expensive. The reward is defined as $R(s_{n})=w\cdot\mathbb{I}(\text{success})$. In addition, we augment the above reward with the process reward model (PRM) using Qwen2.5-Math-PRM-7B (Zhang et al., 2025) to evaluate the intermediate reasoning steps: $R(s_{n})=w\cdot\mathbb{I}(\text{success})+\lambda\sum_{t=1}^{n}\text{PRM}(a_{t}|s_{t-1},g)$.